OpenStudy (anonymous):

What is the axis of symmetry for f(x) = −5x2 − 20x − 10

OpenStudy (campbell_st):

for a parabola \[y = ax^2 + bx + c\] the axis of symmetry is \[x = \frac{-b}{2 \times a}\] in your problem you have a = -5 and b = -20 just substitute and get a value for x, which is the equation of the axis of symmetry

OpenStudy (tkhunny):

If you are really ambitious, you can "Complete the Square". Same result, of course.

OpenStudy (crabbyoldgamer):

You could also set the derivative = to 0, then solve for x.